What rate of interest compounded continuously is needed for an investment of $500 to grow to $900 in 10 years?
Answer by lwsshak3(11628)

What rate of interest compounded continuously is needed for an investment of $500 to grow to $900 in 10 years?
**
Compound Interest Formula For Continuous Compounding: A=Pe^rt, P=initial investment, r=annual rate of interest, A=amount after t years.
For given problem:
900=500e^10r
900/500=e^10r=9/5
take log of both sides
10r*lne=ln (9/5)
log of base=1
10r=ln(9/5)=.5878
r=.5878/10=.05878
ans:
Annual rate of interest compounded continuously needed=5.9%